OpenStudy (anonymous):

write an equation of the line which is parallel to y=1/2x=4 and passes through the point(6,-4)

OpenStudy (anonymous):

do mean y=1/2x+4 ?

OpenStudy (anonymous):

yes i meant that, my bad can u pls help me?

OpenStudy (anonymous):

yes i can :)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

\[y=(1/2)x-7\]

OpenStudy (anonymous):

you can also use this website in the future http://www.algebra.com/algebra/homework/Linear-equations/equation-parallel-or-perpendicular.solver

OpenStudy (anonymous):

okaii thanks a lot but how did u find the -7?

OpenStudy (anonymous):

you have to use the point-slope formula to figure out the equation of the line so y - y1=m(x-x1) so y+4=(1/2)*(x-6) then distribute the 1/2 then get y on its own. when you reduce fractions, you will get the -7

OpenStudy (anonymous):

y - y1=m(x-x1) y+4=(1/2)*(x-6) y+4=(1/2)*x-6/2 y=1/2x -6/2 -4 or y=1/2x -6/2 -8/2 y=1/2x-7

OpenStudy (anonymous):

omg thank you soo much:D

OpenStudy (anonymous):

No problem :)